OpenStudy (anonymous):

simplify the square root of 98x^5

mathslover (mathslover):

can i write 98 as 49 * 2 ?

OpenStudy (anonymous):

\[\sqrt{98x^5}= \sqrt{2*49*x^4*x}\]

OpenStudy (anonymous):

uh i guess i think the answer is 2x^2 √7x is that right?

OpenStudy (anonymous):

how did you factor out a 2???

OpenStudy (anonymous):

\[2x ^{2}\sqrt{7x}\] is that right ??

OpenStudy (anonymous):

it is incorrect... where did you get the 2 from?

OpenStudy (anonymous):

and where did you get the 7 from

OpenStudy (anonymous):

um i guess i did it wrong i have four choices that was the first one i tried

OpenStudy (anonymous):

all of the choices involve a 7 and a 2

OpenStudy (anonymous):

... ok then what is the square root of 49?

OpenStudy (anonymous):

I think it 7

OpenStudy (anonymous):

good, that means you can factor out a 7 what is the square root of x^4?

OpenStudy (anonymous):

could it be \[7x ^{2}\sqrt{2x}\]

OpenStudy (anonymous):

\[\sqrt{98x^5}=\sqrt{49*2*x^4*x}=7\sqrt{2*x^4*x}\]

OpenStudy (anonymous):

\[\sqrt{98x^5}=\sqrt{49*2*x^4*x}=7\sqrt{2*x^4*x}=7x^2 \sqrt{2x}\]

OpenStudy (anonymous):

Thank you so much for your help i really really do appreciate it sorry im not really good at math